OpenStudy (jpsmarinho):

lim [(sin(x^2+2)-sin(x+2))/x] when x goes to 0. I should use sin(a+b) = sinacosb + sinbcosa?

OpenStudy (anonymous):

Are you allowed to use L'hopital's rule?

OpenStudy (anonymous):

\[\Large \lim_{x \rightarrow 0} \frac{\sin(x^2+2)-\sin(x+2)}{x}\]

OpenStudy (anonymous):

If you know L'hopital's rule, this problem is pretty simple, but there's a possibility you haven't been taught that yet. Let me know if you're allowed.

OpenStudy (jpsmarinho):

Yeah, I'm allowed to use L'Hôspital rule, but doing in "pure algebra" is nice to learn, I think. ^^

OpenStudy (jpsmarinho):

How do you wrote the lim in a bigger font

OpenStudy (jpsmarinho):

?

OpenStudy (anonymous):

Use the equation editor. and I typed: Large lim_{x rightarrow 0} frac{sin(x^2+2)-sin(x+2)}{x} Large makes it bigger, and typing Large makes it larger than typing large. The only other part that might be hard to figure out is: frac{numerator}{denominator} makes a fraction

OpenStudy (jpsmarinho):

Thanks!

OpenStudy (anonymous):

You're welcome. Okay, so here's how to use L'hopital's on this. First, make sure that you're allowed to use the rule. Is the function of indeterminate form? Well, if we put in x=0, we get lim = 0/0, so yes, we can use L'hopital's rule. Simply take the derivative of top and bottom: \[\Large = \lim_{x \rightarrow 0} \frac{2x*\cos(x^2+2)-\cos(x+2)}{1} = -\cos(2)\]

OpenStudy (anonymous):

Now, as for solving it algebraically, I think you're on the right track... but that's a good amount of work, and frankly, I don't want to do it, since I can just L'hopital this bad boy right in the throat.

OpenStudy (jpsmarinho):

Ah it's true that is much easier use L'Hôspital in this case. Thanks!! A doubt : I'm brazilian, and I don't understand very nice this: "right in the throat." is this only a expression that means accept something? I'm curious.

OpenStudy (anonymous):

lol sorry about that. That is not a common phrase at all, and hardly worth explaining.

OpenStudy (anonymous):

I just meant to express that it is kind of an easy solution to a difficult problem.

OpenStudy (jpsmarinho):

Ahh right ! Rsrs thanks again!